OpenStudy (anonymous):

When viewing Big Ben in London, you decide that you want to visit the Jubilee Gardens. Instead of traveling by car, you take a boat ride to the Jubilee Gardens. Using the map below, calculate the distance you traveled from Big Ben to the Jubilee Gardens by boat. Show all work and explain each step of your calculations.

OpenStudy (pratyush5):

map ?

OpenStudy (anonymous):

Sorry, I don't know how to include the map but if traveling by road you go 7 miles east and 5 miles north. So, the distance by boat would make a triangle.

OpenStudy (jim766):

 this creates a right triangle, you can use Pythagorean thermo a^2 + b^2 = c^2 c is the side you are trying to find, I assume

OpenStudy (jim766):

5^2 + 6^2 = c^2

OpenStudy (jim766):

25 + 36 = c^2

OpenStudy (jim766):

sorry I mistyped 5^2 + 7^2 = c^2

OpenStudy (jim766):

25 + 49 = c^2 74 = c^2 now to solve, take sq root of both sides \[\sqrt{74}= c\] \[8.6\approx c \]

OpenStudy (anonymous):

thank you! I've got it now!

OpenStudy (jim766):

yw